Flying with the wind, a plane flew 1000 miles in 4 hours. Flying against the wind, the plane could fly only 500 miles in the same amount of time. Find the rate of the plane in calm air and the rate of the wind.
Answer by checkley77(12844)

d=rt
1000=4r
r=1000/4
r=250 mph. plane speed with the wind.
500=4r
r=500/4
r=125 mph. plane speed against the wind.
(250-125)/2=125/2=62.5 mph. was the wind speed.
250-62.5=187.5 was the speed of the plane in calm air.
125+62.5=187.5 ditto.